# Codec Does Matter: Exploring the Semantic Shortcoming of Codec for Audio Language Model

### v2

@@ Methods
Table 5: Performance of semantic representation on the ARCH benchmark. The table shows the performance of various models across different domains. ESC-50 (Piczak 2015), US8K (Salamon, Jacoby, and Bello 2014), FSD50K (Fonseca et al. 2022), and VIVAE (Holz, Larrouy-Maestri, and Poeppel 2022) represent performance on Acoustic Events. FMA (Defferrard et al. 2017), MTT (Law et al. 2009), IRMAS (Bosch et al. 2012), and MS-DB (Bittner et al. 2014) indicate performance in the Music domain. RAVDESS (Livingstone and Russo 2018), AudioMNIST (Becker et al. 2024), SLURP (Bastianelli et al. 2020), and EMOVO (Costantini et al. 2014) reflect performance in the Speech domain. Higher values indicate better performance across these tasks.
Training Details: Still, GPT-2-medium (approximately 300M parameters) are adopted for conditional text-to-sound tasks, where the condition embedding is extracted from text captions using LAION-CLAP (Wu et al. 2023c) and linearly projected from 512 dimensions to 1024 dimensions for GPT input. The training data consists of approximately 400 hours of audio content sourced from the AudioCaps dataset (Kim et al. 2019) and the AudioSet SL subset from the WavsCaps dataset (Mei et al. 2024). All audio samples are uniformly resampled to a 16kHz sampling rate. The first four tokens from the VQ layers are preprocessed and flattened to configure the GPT model’s block size to 2000, corresponding to a processing rate of 50Hz. The training process spans 80,000 steps on four NVIDIA 4090 GPUs, with a batch size of 8 and a learning rate of 3e-4. A warmup phase of 2000 steps is employed to optimize the training process.
Evaluation Metrics: following (Huang et al. 2023) and (Liu et al. 2023), we calculate Frechet Distance (FD), Inception Score (IS), Frechet Audio Distance (FAD) for text-to-audio generation. In addition, CLAP score (Huang et al. 2023) is used to evaluate the correspondence between the generated audio and the text prompt.
@@
Analysing the Effect of Codec
We hypothesize that the enhanced audio generation capabilities of the audio LLMs are attributed to the improved semantic understanding facilitated by the X-Codec. To validate this hypothesis, we employ the ARCH benchmark (La Quatra et al. 2024) to evaluate the audio semantic understanding, and the benchmark is a comprehensive framework specifically designed to evaluate automatic recognition learning methods across a diverse range of audio classification domains, including acoustic events, music, and speech. The results from this benchmark are shown in Table 5.
@@
Limitation
While our method significantly enhances the performance of codecs for LLMs by integrating semantic information, it does come with certain trade-offs. According to the principle of ”no free lunch,” improving one aspect of a system often involves compromises in others. In the case of our enhanced codecs, the primary limitation lies in their potential impact on the original functionality of codecs, which is compression for information transmission. The introduction of a semantic extraction layer adds additional computational overhead, potentially increasing the time required for processing. This can affect the efficiency of the codec when used in applications where rapid data compression and transmission are critical. Consequently, while our approach offers substantial benefits for semantic understanding and audio processing, it may not be as effective in contexts where high-speed data compression is paramount.

### v3

@@ 4.2.3 Text-to-Sound
Table 5: Performance of semantic representation on the ARCH benchmark. The table shows the performance of various models across different domains. ESC-50 [36], US8K [37], FSD50K [38], and VIVAE [39] represent performance on Acoustic Events. FMA [40], MTT [41], IRMAS [42], and MS-DB [43] indicate performance in the Music domain. RAVDESS [44], AudioMNIST [45], SLURP [46], and EMOVO [47] reflect performance in the Speech domain. Higher values indicate better performance across these tasks.
Training Details: Still, GPT-2-medium (approximately 300M parameters) are adopted for conditional text-to-sound tasks, where the condition embedding is extracted from text captions using LAION-CLAP [48] and linearly projected from 512 dimensions to 1024 dimensions for GPT input. The training data consists of approximately 400 hours of audio content sourced from the AudioCaps dataset [49] and the AudioSet SL subset from the WavsCaps dataset [50]. All audio samples are uniformly resampled to a 16kHz sampling rate. The first four tokens from the VQ layers are preprocessed and flattened to configure the GPT model’s block size to 2000, corresponding to a processing rate of 50Hz. The training process spans 80,000 steps on four NVIDIA 4090 GPUs, with a batch size of 8 and a learning rate of 3e-4. A warmup phase of 2000 steps is employed to optimize the training process.
Evaluation Metrics: following [51] and [52], we calculate Frechet Distance (FD), Inception Score (IS), Frechet Audio Distance (FAD) for text-to-audio generation. In addition, CLAP score [51] is used to evaluate the correspondence between the generated audio and the text prompt.
@@
4.2.4 Analysing the Effect of Codec
We hypothesize that the enhanced audio generation capabilities of the audio LLMs are attributed to the improved semantic understanding facilitated by the X-Codec. To validate this hypothesis, we employ the ARCH benchmark [53] to evaluate the audio semantic understanding, and the benchmark is a comprehensive framework specifically designed to evaluate automatic recognition learning methods across a diverse range of audio classification domains, including acoustic events, music, and speech. The results from this benchmark are shown in Table 5.
@@ 4.2.4 Analysing the Effect of Codec
4.3 Limitation
While our method significantly enhances the performance of codecs for LLMs by integrating semantic information, it does come with certain trade-offs. According to the principle of "no free lunch," improving one aspect of a system often involves compromises in others. In the case of our enhanced codecs, the primary limitation lies in their potential impact on the original functionality of codecs, which is compression for information transmission. The introduction of a semantic extraction layer adds additional computational overhead, potentially increasing the time required for processing. This can affect the efficiency of the codec when used in applications where rapid data compression and transmission are critical. Consequently, while our approach offers substantial benefits for semantic understanding and audio processing, it may not be as effective in contexts where high-speed data compression is paramount.